You are planning to retire in twenty years. You'll live ten years after retirement. You want to be able to draw out of your savings at the rate of $10,000 per year. How much would you have to pay in equal annual deposits until retirement to meet your objectives? Assume interest remains at 9% ?

thank you
1
Expert's answer
2015-01-09T08:56:06-0500
FV = 10*$10,000, r = 9%, n = 20
In this case we use the formula FV = PV*(1+r)^n
PV*(1.09^1 + 1.09^2 + 1.09^3 + 1.09^4 + 1.09^5 + 1.09^6 + 1.09^7 + 1.09^8 + 1.09^9 + 1.09^10 + 1.09^11 + 1.09^12 + 1.09^13 + 1.09^14 + 1.09^15 + 1.09^16 + 1.09^17 + 1.09^18 + 1.09^19 + 1.09^20) = 10,000*10
PV =100,000/(1.09^1 + 1.09^2 + 1.09^3 + 1.09^4 + 1.09^5 + 1.09^6 + 1.09^7 + 1.09^8 + 1.09^9 + 1.09^10 + 1.09^11 + 1.09^12 + 1.09^13 + 1.09^14 + 1.09^15 + 1.09^16 + 1.09^17 + 1.09^18 + 1.09^19 + 1.09^20) = 1793.26